### v1

# MMASD: A Multimodal Dataset for Autism Intervention Analysis

Abstract.
Autism spectrum disorder (ASD) is a developmental disorder characterized by significant social communication impairments and difficulties perceiving and presenting communication cues. Machine learning techniques have been broadly adopted to facilitate autism studies and assessments. However, computational models are primarily concentrated on specific analysis and validated on private datasets in the autism community, which limits comparisons across models due to privacy-preserving data sharing complications. This work presents a novel privacy-preserving open-source dataset, MMASD as a MultiModal ASD benchmark dataset, collected from play therapy interventions of children with Autism. MMASD includes data from 32 children with ASD, and 1,315 data samples segmented from over 100 hours of intervention recordings. To promote public access, each data sample consists of four privacy-preserving modalities of data; some of which are derived from original videos: (1) optical flow, (2) 2D skeleton, (3) 3D skeleton, and (4) clinician ASD evaluation scores of children, e.g., ADOS scores. MMASD aims to assist researchers and therapists in understanding children’s cognitive status, monitoring their progress during therapy, and customizing the treatment plan accordingly. It also has inspiration for downstream tasks such as action quality assessment and interpersonal synchrony estimation. MMASD dataset can be easily accessed at https://github.com/Li-Jicheng/MMASD-A-Multimodal-Dataset-for-Autism-Intervention-Analysis.
multimodal dataset, machine learning, deep learning, autism spectrum disorder, human activity recognition
journalyear: 2023copyright: acmcopyrightconference: INTERNATIONAL CONFERENCE ON MULTIMODAL INTERACTION; October 9–13, 2023; Paris, Francebooktitle: INTERNATIONAL CONFERENCE ON MULTIMODAL INTERACTION (ICMI ’23), October 9–13, 2023, Paris, Franceprice: 15.00ccs: Social and professional topics People with disabilitiesccs: Human-centered computingccs: Computing methodologies Activity recognition and understandingccs: Computing methodologies Machine learning
Figure 1. MMASD provides multiple multimodal privacy-preserving features derived from original videos via ROMP (Sun et al., 2021), OpenPose (Cao et al., 2019) and Lucas-Kanade (Lucas and Kanade, 1981), including optical flow, 2D/3D skeleton and clinician evaluation results. The dataset can be accessed publicly and investigate research questions centered on social, and behavioral interactions of children with ASD in playful, group activities.
taser
Figure 1. MMASD provides multiple multimodal privacy-preserving features derived from original videos via ROMP (Sun et al., 2021), OpenPose (Cao et al., 2019) and Lucas-Kanade (Lucas and Kanade, 1981), including optical flow, 2D/3D skeleton and clinician evaluation results. The dataset can be accessed publicly and investigate research questions centered on social, and behavioral interactions of children with ASD in playful, group activities.
1. Introduction
Autism spectrum disorder (ASD) is a neurodevelopmental disorder characterized by significant impairments in social communication, as well as difficulties in perceiving and expressing communication cues. Approximately 1 in 54 children are on the spectrum in the United States, resulting in over 1 million affected individuals nationwide (Baio et al., 2018). Primary treatment of ASD includes behavioral and psycho-social interventions accompanied by prescribed medication. Behavioral and psycho-social interventions facilitate social and communication development, while medication helps control associated symptoms and comorbid problems (Billing et al., 2020). Specifically, psycho-social interventions are diverse in content and can vary in curriculum, structure, discipline, and theme. Prevailing therapeutic interventions include applied behavior analysis and robot-assisted therapy, both of which can provide valuable data for analyzing children’s mental development and developing individualized treatment plans.
Various studies have widely adopted machine learning techniques to facilitate autism research (Li et al., 2022a, b; Guo et al., 2023; Marinoiu et al., 2018; Pandey et al., 2020; Chen and Zhao, 2019; de Belen et al., 2020). Compared to traditional methods that rely heavily on human expertise and experience, machine learning approaches can help reduce the need for human labor and associated costs while still achieving decent performance. The autism community has benefited from machine learning techniques in many areas, including but not limited to autism diagnosis (Chen and Zhao, 2019; Wall et al., 2012), emotion recognition (Li et al., 2021b; Marinoiu et al., 2018), and movement pattern assessment (Li et al., 2022a, b, 2021a; Pandey et al., 2020).
In machine learning research, it is widely accepted to follow a research pipeline that involves developing, applying, and comparing models across multiple benchmark datasets to ensure fair comparisons in performance. However, in the autism community, commonly recognized benchmarks, especially for behavior analysis and activity understanding, are limited due to privacy concerns. Typically, models are validated on a private dataset, and data-sharing roadblocks can restrict the comparison between models. In this sense, the availability of publicly accessible datasets is a crucial first step for the autism community since it allows cutting-edge machine-learning techniques to be trained and validated on ASD datasets. Although some studies have already been conducted, there is still a significant lack of publicly available, multimodal datasets that can be used to analyze the full-body movements of children during therapeutic interventions.
To overcome the data-sharing roadblock, in this paper, we proposed a publicly available multimodal dataset, MMASD [^1] MMASD can be accessed at https://github.com/Li-Jicheng/MMASD-A-Multimodal-Dataset-for-Autism-Intervention-Analysis., collected from physical therapy interventions of children with autism. MM-ASD maintains privacy while retaining essential movement features by providing optical flow, 2D and 3D skeletons that are derived from the original videos, thereby avoiding the exposure of sensitive and identifiable raw video footage. Additionally, it includes clinician evaluation results of each child such as ADOS-2 and motor function scores. We also provide the intervention activity class labels for overall scene understanding. Overall, MMASD can be used to help therapists and researchers to understand children’s cognitive status, track development progress in therapy, and guide the treatment plan accordingly. It also provides inspiration for downstream tasks, such as activity recognition (Pandey et al., 2020), action quality assessment (Li et al., 2021a), and interpersonal synchrony estimation (Li et al., 2022b). In contrast to current datasets, e.g., listed in Table  1, our dataset stands out for the following reasons:
  • It is a publicly accessible benchmark dataset for movement and behavior analysis during therapeutic interventions featured by diverse scenes, group activities, and participants.
  • It includes multimodal features such as optical flow, 2D/3D skeletons, demographic data, and ADOS score rating. These features provide privacy-preserving approaches to maintain critical and full-body motion information.
  • Each scene depicts the same activity performed by a child and one or more therapists, providing an in-place template for comparing typically developing individuals with children with ASD.
The rest of the paper is organized as follows. Further background on relevant autism datasets is presented in Section 2, followed by a presentation of data collection approach in Section 3. Details of the dataset, including statistics, data processing and annotation, are provided in Section 4. Finally, discussion and current limitations of MMASD are described in Section 5 and conclusion in Section 6.
Table 1. A comparison of related benchmarks and our dataset focused on behavior and movement analysis of children with ASD, including the target population, research focus, data modalities, computational models, data sample, and availability.
Dataset
Target
Population
Research Focus
Data Modalities
Computational Models
$N$ of Data Samples
(Data length)*
Availability
Billing et al. (Billing et al., 2020)
61 ASD
Behavior analysis
Body motion, head pose, eye gaze
SVM, face detector (Viola and Jones, 2004), Microsoft Kinect SDK
3121 sessions (306 h)
Public
Rajagopalan et al. (Rajagopalan et al., 2013)
Self-stimulatory behaviors detection
Video, audio
Space time interest point (STIP) (Laptev, 2005)
75 videos (90 s/video)
Public
Rehg et al. (Rehg et al., 2013)
121 total
Behavior analysis
Video, audio, physiological data
SVM, Omron OKAO vision library
160 sessions (3–5 m/video)
On request
DE-ENIGMA (Riva et al., 2020)
128 ASD
Behavior analysis
Facial landmark, body postures, video, audio
Computer vision, Microsoft Kinect SDK
50 annotated recordings (154 h)
On request
Zunino et al. (Zunino et al., 2018)
20 ASD, 20  TD
Grasping actions
Hand & arm trajectories
LSTM, CNN, Histogram of optical flow (HOG), VLAD (Jégou et al., 2011)
1837 videos (83 frames/video)
On request
Pandey et al. (Pandey et al., 2020)
37 ASD
Behavior analysis
Video, 2D pose, optical flow
Guided weak supervision, Temporal segment networks, Inflated 3D CNN
1481 video clips
On request
Del Coco et al. (Del Coco et al., 2017)
8 ASD
Behavior analysis
Facial landmark, gaze, head pose
Constrained Local Neural Fields (Baltrusaitis et al., 2013)
6 videos (joint activities), 4 videos (imitative play)
Private
Dawson et al. (Dawson et al., 2018)
22 ASD, 82 TD
Phenotyping, Head movement
Head pose, facial landmark
Model-based object pose, Computer vision analysis (CVA) (la Torre et al., 2015)
10 videos (1–3 s/video, 49 facial landmarks)
Private
Martin et al. (Martin et al., 2018)
21 ASD, 21 TD
Head movement
Head pose, facial landmark
Computer-vision head tracking (Zface) (Jeni et al., 2015)
252 videos (6 m/video)
Private
MMASD (Ours)
32 ASD
Movement & Behavior analysis
Optical flow, 2D pose, 3D pose, ADOS score
ROMP (Sun et al., 2021), OpenPose (Cao et al., 2019), Lucas-Kanade (Lucas and Kanade, 1981)
1,315 videos (7 s & 186 frames/video)
Public
*Certain works provided information on the duration of individual videos, whereas others presented the overall length of the dataset.
2. Related Work
ASD is characterized by atypical movement patterns, such as repetitive movements, clumsiness, and difficulties with coordination. A deeper understanding of these movement patterns and their association with ASD can aid therapists, clinicians, and researchers in developing more effective interventions and therapies. To this end, several datasets have been developed for movement analysis of children with ASD. These datasets typically involve collecting motion capture or sensor data from children while they perform various activities or specific tasks. The collected data is then analyzed to identify differences and patterns in movement between children with ASD and typically developing children. In this section, we present an overview of existing datasets that focus on movement and behavior analysis of children with ASD.
Movement analysis with humanoid robot interactions datasets
Marinoiu et al. (Marinoiu et al., 2018) presented a dataset and system that use a humanoid robot to interact with children with ASD and monitor their body movements, facial expressions, and emotional states. The results show that children significantly improved their ability to recognize emotions, maintain eye contact, and respond appropriately to social cues. They identified several effective techniques for detecting and analyzing the children’s emotional and behavioral responses, such as analyzing the frequency and duration of specific behaviors. Billing et al. (Billing et al., 2020) proposed a dataset of behavioral data recorded from 61 children with ASD during a large-scale evaluation of robot-enhanced therapy. The dataset comprises sessions where children interacted with a robot under the guidance of a therapist and sessions where children interacted one-on-one with a therapist. For each session, they used three RGB and two RGBD (Kinect) cameras to provide detailed information, including body motion, head position and orientation, and eye gaze of children’s behavior during therapy.
Another dataset related to humanoid robot interactions with autistic children is DE-ENIGMA (Riva et al., 2020), which includes using a multimodal human-robot interaction system to teach and expand social imagination in children with ASD. The DE-ENIGMA dataset comprises behavioral features such as facial mapping coordinates, visual and auditory, and facilitates communication and social interaction between the children and the robot. The authors indicated that the DE-ENIGMA could be used as an effective tool for teaching and expanding social imagination in children with ASD. They also suggest that the usage of a multimodal human-robot interaction could be a promising approach for developing interventions for children with ASD that aim to improve their social skills and promote better social integration.
Eye movement and vocalization datasets
Duan et al. (Duan et al., 2019) introduced a dataset of eye movement collected from children with ASD. The dataset includes 300 natural scene images and eye movement data from 14 children with ASD and 14 healthy individuals. It was created to facilitate research on the relationship between eye movements and ASD, with the goal of designing specialized visual attention models. Baird et al. (Baird et al., 2017) introduced a dataset of vocalization recordings from children with ASD. They also evaluated classification approaches from the spectrogram of autistic speech instances. Their results suggest that automatic classification systems could be used as a tool for aiding in the diagnosis and monitoring of ASD in children.
Behavior analysis datasets
For action recognition dataset of children with ASD, Pandey et al. (Pandey et al., 2020) proposed a dataset of video recording actions and a technique to automate the response of video recording scenes for human action recognition. They evaluated their technique on two skill assessments with autism datasets and a real-world dataset of 37 children with ASD. Rehg et al. (Rehg et al., 2013) introduced a publicly available dataset including over 160 sessions of child-adult interactions. They discussed the use of computer vision and machine learning techniques to analyze and understand children’s social behavior in different contexts. They also identified technical challenges in analyzing various social behaviors, such as eye contact, smiling, and discrete behaviors. Rajagopalan et al. (Rajagopalan et al., 2013) explored the use of computer vision techniques to identify self-stimulatory behaviors in children with ASD. They also presented a self-stimulatory behavior dataset (SSBD) to assess the behaviors from video records of children with ASD in uncontrolled natural settings. The dataset comprises 75 videos grouped into three categories: arm flapping, head banging, and spinning behaviors.
Comparison to MMASD
In Table 1, we compare our proposed dataset with related benchmarks. Overall, MMASD features diverse themes and scenes, capturing full-body movements with multimodal features. In contrast, some works focused specifically on upper-body movements (Zunino et al., 2018; Del Coco et al., 2017; Dawson et al., 2018; Martin et al., 2018). MMASD also provides critical privacy-preserving features to represent body movements making it publicly accessible, while some works were conducted on raw videos that are either private or accessible only upon request (Billing et al., 2020; Rehg et al., 2013; Riva et al., 2020; Del Coco et al., 2017; Dawson et al., 2018; Martin et al., 2018). Additionally, it is collected from therapeutic interventions, reflecting participants’ motor ability and providing valuable insights for treatment guidance.
3. Method
In the following sections, we describe the participants, procedure, and experimental settings of our proposed dataset.
3.1. Participants
We recruited 32 children (27 males and 5 females) with ASD from different races (Caucasian, African American, Asian, and Hispanic) and backgrounds through fliers posted online and onsite in local schools, services, and self-advocacy groups. Prior to enrollment, children were screened using the Social Communication Questionnaire (Srinivasan et al., 2016), and their eligibility was determined by the Autism Diagnostic Observation Schedule-2 (ADOS-2) (Lord et al., 2000; Srinivasan et al., 2015b) as well as clinical judgment. All the children were between 5 and 12 years old. Written parental consent was obtained from the Institutional Review Board of the university before enrollment. The Vineland Adaptive Behavior Scales (Sparrow and Cicchetti, 1989) were used to assess the children’s adaptive functioning levels. In general, 82% of the participating children had delays in the Adaptive Behavior Composite. Specifically, 70% of them experienced communication delays, 80% had difficulties with daily living skills, and 82% had delays in socialization
3.2. Procedure
The study was conducted over ten weeks, with the pre-test and post-test being conducted during the first and last weeks of the study, respectively. Each training session was scheduled four times per week and lasted approximately 45 minutes. During the intervention, the trainer and adult model interacted with the child within a triadic context, with the adult model acting as the child’s confederate and participating in all activities with the child. This triadic setting (child, trainer, and model) provided numerous opportunities for promoting social and fine motor skills such as eye contact, body gesturing and balancing, coordination, and interpersonal synchrony during joint action games.
All expert trainers and models involved were either physical therapists or physical therapy/kinesiology graduate students who had received significant pediatric training prior to their participation. The trainers and models were unknown to the children before the study. In addition to the expert training sessions, we also encouraged parents to provide two additional weekly sessions involving similar activities to promote practice. Parents were provided with essential instruction manuals, supplies, and in-person training beforehand. All training sessions were videotaped with the parents’ consent and notification to the children, and the training diary was compiled by parents in collaboration with expert trainers. The general pipeline of training sessions had a standard procedure despite some unique activities across different themes. A welcoming and debriefing phase was present at the beginning and end of the data collection to help children warm up and get ready for the intervention, as well as to facilitate the subsequent data processing stage by providing time labels that indicate the segments to investigate.
3.3. Experiment Settings
All videos were recorded in a house environment with the camera pointed toward the children. Models (trainers/therapists and trained adults) interacted with the child within a triadic context, and the adult model was the child’s confederate and practiced all activities with the child. Different tools were introduced to facilitate the training process depending on the theme of the intervention, for example, instruments and robots. Selected scenes in different themes of our proposed MMASD dataset are shown in Figure 2.
Figure 2. Sample scenes depicting various themes and activity classes present in the MMASD dataset.
4. MMASD
MMASD includes 32 children diagnosed with autism of different levels. It covers three unique themes:
  • Robot: children followed a robot and engaged in body movements.
  • Rhythm: children and therapists played musical instruments or sang together as a form of therapy.
  • Yoga: children participated in yoga exercises led by experts. These exercises included body stretching, twisting, balancing, and other activities.
Overall, MMASD comprises 1,315 video clips that have been meticulously gathered from intervention video recordings spanning more than 108 hours. It consists of 244,679 frames with an average duration of 7.15 seconds. The average data length in MMASD is $7.0\pm 3.4$ seconds ($186.1\pm 92.9$ frames), with dimensions ranging from $320\times 240$ to $1920\times 1080$. Table 2 presents statistical information on MMASD. Depending on the activity conducted during the intervention, we further categorized all data into eleven activity classes as described in Table 3. Each activity class falls into a unique theme, as shown in Figure 2. MMASD also reports demographic and autism evaluation scores of all participating children, including date of birth, ADOS-2 score (social affect & restricted and repetitive behavior), motor functioning score, and severity of autism.
Table 2. Statistics of our proposed MMASD dataset.
Description Value
Number of data samples 1,315
Number of frames 244,679
Number of activity classes 11
Average video length (seconds) $7.0\pm 3.4$
Average number of frames $186.1\pm 92.9$
Resolution $320\times 240$ $\sim$ $1920\times 1080$
FPS 25 $\sim$ 30
Table 3. Description and distribution of all 11 activity classes in MMASD.
Activity Class
Activity Description
Count
Arm swing
The participant raises their left and right arm in succession while maintaining an upright posture.
105
Body swing
The participant swings their body left and right while stretching out both hands, one behind the other.
119
Chest expansion
The participant gradually opens and closes their chest.
114
Drumming
The participant plays either the snare or Tubano drum with one or both hands.
168
Frog pose
The participant widens their knees as far as possible, places their feet with the big toes touching behind the buttocks, and assumes a shape like that of a frog in the kneeling position.
113
Maracas forward shaking
The participant shakes maracas back and forth, an instrument commonly appearing in Caribbean and Latin music.
103
Maracas shaking
The participant shakes maracas left and right in front of their chest.
130
Sing and clap
The participant sits on the ground while simultaneously singing and clapping, typically done at the start or end of an intervention.
113
Squat
The participant repeatedly performs a crouching stance with their knees bent.
101
Tree pose
The participant balances on one leg and places the sole of the other foot on the inner thigh, calf, or ankle of the standing leg in tree pose.
129
Twist pose
The participant sits with their legs crossed and twists their torso to one side, keeping their lower body stable and grounded.
120
4.1. Data processing
From original video recordings, we manually find out the start and end time stamps of a specific activity. Then we segmented the video into clips and categorized them by activity class. Clips shorter than three seconds were discarded. We also discarded noisy data due to video quality, lighting conditions, and body occlusion. Besides all the eleven activities in MMASD, there were some other activities with fewer examples. To ensure a balanced data distribution, we excluded all inadequate classes.
4.2. Data Annotation
We have four annotators that are well-trained in intervention understanding. The annotators had a comprehensive understanding of interventions and a cross-disciplined background in computer science and physical therapy. We exclusively assigned one activity class label to each video. Each annotator completed data annotation independently, and the final class label was determined by majority voting. However, original videos cannot be publicly shared due to privacy concerns. Therefore, we create data samples from every video clip by extracting selected features from the original scenes, including (1) optical flow, (2) 2D skeleton, and (3) 3D skeleton, respectively. All these features can maintain critical body movements while preserving privacy. Section 4.3 explained all selected features in detail.
In addition to the motion-related features mentioned above, we also reported clinician evaluation results such as ADOS-2 score, motor functioning score, and severity of autism for each participating child. ADOS-2 is a standardized assessment tool used to evaluate individuals suspected of having ASD. It is used in conjunction with other diagnostic information to help clinicians determine whether an individual meets the criteria for an ASD diagnosis. ADOS-2 includes several modules, each designed for individuals of different ages and language abilities and includes a series of activities and tasks that are used to observe key features of ASD, such as social communication skills and repetitive behaviors. The ADOS-2 scores are based on an individual’s performance during activities and tasks specific to their module and can range from 0 to 10 or higher depending on the algorithm used, with higher scores indicating more severe ASD symptoms. Moreover, we reported the ADOS comparison score, a continuous metric ranging from 1 to 10 that describes the severity of a child’s autism symptoms compared to children with ASD of similar age and language levels (Gotham et al., 2009). Low comparison scores are indicative of minimal evidence of autism symptoms, whereas high scores are indicative of severe autism symptoms.
The contrast between the ADOS-2 score and the ADOS comparison score is worth mentioning. The ADOS-2 score reflects an individual’s raw score on the ADOS-2 assessment tool, while the ADOS comparison score is a statistical measure that compares an individual’s performance to others of the same age and language level. The motor functioning score refers to an assessment of an individual’s motor skills and abilities and is evaluated based on the children’s level of independence in daily living skills (Srinivasan et al., 2015a, 2016). It is on a scale of 1 to 3, while $1,2,3$ represent low functioning (needing significant support), medium functioning (needing moderate support), and high functioning (needing less support), respectively. Finally, the severity of autism is determined by a comprehensive assessment that includes both the ADOS and motor function evaluation.
4.3. Feature Extraction
In order to preserve critical details of movement while avoiding any infringement of privacy, we derived the subsequent features from the initial footage.
(1) Optical flow An optical flow is commonly referred to as the apparent motion of individual pixels between two consecutive frames on the image plane. Optical flow derived from raw videos (see Figure 1) can provide a concise description of both the region and velocity of a motion without exposing an individual’s identity (Carreira and Zisserman, 2017; Feichtenhofer et al., 2016; Gao et al., 2020; Pandey et al., 2020).
(2) 2D Skeleton Skeleton data has an edge over RGB representations because it solely comprises the 2D positions of the human joints, which offer highly conceptual and context-independent data. This allows models to concentrate on the resilient aspects of body movements. 2D skeleton data has been widely applied to tasks relating to human behavior understanding, such as action recognition (Zheng et al., 2021; Yan et al., 2018), action quality assessment (Li et al., 2022b, a) and beyond. An optimal way to acquire skeleton data is through the use of wearable devices and sensors that are affixed to the human body. However, in the context of autism research, it poses a substantial challenge as children may feel overwhelmed wearing these devices and experience anxiety. As a result, the skeleton data extraction process is carried out by pre-trained pose detectors based on deep neural networks.
(3) 3D Skeleton Similar to 2D skeleton data, 3D skeletons instead represent each key joint with a 3D coordination, introducing an additional depth dimension. Since all the data was collected using a single RGB camera, we also completed this process with the help of deep neural networks.
The technical details and tools utilized for feature extraction can be found in Section 4.5.
4.4. Data Format
Suppose the original video clip includes $N$ participants (child, trainer, and assistant) is composed of $L$ frames, and the height and width of each frame are $H$ and $W$, respectively. As discussed above, each data sample consists of four distinct components, with data dimension demonstrated in braces:
  • Optical flow $(L-1,H,W)$: saved as npy files (Harris et al., 2020).
  • 2D skeleton $(L,N,17,2)$: 2D coordinates of 17 key joints, following COCO (Coco and Dale, 2014) format, saved as JSON files.
  • 3D skeleton $(L,N,24,3)$: 3D coordinates of 24 key joints, following ROMP (Sun et al., 2021) format, saved as npz files.
  • Demographic and clinical evaluation for ASD (9,)(9,)( 9 , ): including nine attributes such as participant ID, date of birth, chronological age, ADOS-2 module, social affect score, restricted and repetitive behavior score, ADOS comparison score, motor functioning score, and severity of autism, saved as CSV files.
4.5. Implementation Details
4.5.1. Optical flow:
The Lucas-Kanade (Lucas and Kanade, 1981) method is used for our study. It is a popular technique used in computer vision to estimate the motion of objects between consecutive frames. The method assumes that the displacement of the image contents between two nearby instants is small and approximately constant within a neighborhood of the point under consideration. By solving the optical flow equation for all pixels within a window centered at the point, the method can estimate the motion of objects in the image sequence. Overall, the Lucas-Kanade optical flow method is an effective and popular technique for estimating motion in various computer vision applications.
4.5.2. 2D skeleton:
The OpenPose method (Cao et al., 2019) is used to extract 2D skeletons from our dataset of human action videos. OpenPose is a powerful tool for body, face, and hand analysis, developed by Carnegie Mellon University, and is based on deep learning techniques. It is a real-time multi-person key-point detection library that can accurately detect the key points of a human body, including joints and body parts, from an image or video feed. Initially, it predicts confidence maps for every body part and subsequently associates them with distinct individuals via Part Affinity Fields. The library is open-source and written in C++ with a Python API, which makes it easy to use and integrate into various computer vision applications.
4.5.3. 3D skeleton:
We utilized the Regression of Multiple 3D People (ROMP) proposed by Sun et al. (Sun et al., 2021), a state-of-the-art technique to estimate the depth and pose of an individual from a single 2D image. The authors proposed a deep learning-based approach that is based on a fully convolutional architecture, which takes an input image and directly predicts the 3D locations of the body joints of the person(s) present in the image. This is achieved by directly estimating multiple differentiable maps from the entire image, which includes a Body Center heatmap and a Mesh Parameter map. 3D body mesh parameter vectors of all individuals can be extracted from these maps using a simple parameter sampling process. These vectors are then fed into the SMPL body model to generate multi-person 3D meshes.
In our study, we employed the code and pre-trained model shared by the authors and used it on our dataset to suit our specific needs. By utilizing this method and applying it to our own data, we obtained 2D and 3D coordinates of key joints of the person(s).
5. Discussion
This section delves into the limitations of MMASD and the challenges we encountered during the dataset preparation process. As the experiments were conducted in real-world settings, we faced common computer vision challenges, including varying video quality, illumination changes, cluttered backgrounds, and pose variations. Notably, in the feature extraction stage, we encountered pose detection failures in challenging scenarios, such as body occlusion and participants moving out of the scene. The intrinsic video quality limitation of MMASD also restricted us from capturing subtle and fine-grained features, such as facial expressions.
Moreover, it is imperative to conduct in-depth investigations into domain-specific challenges. For instance, in standard benchmarks for human activity recognition, typically developing individuals exhibit dominant and continuous actions with similar intensity. However, data on MMASD may not solely contain the target behavior throughout the entire duration, as impromptu actions or distractions may occur during therapy sessions for children. Furthermore, unlike prevailing benchmarks that collect ground truth skeleton data by attaching sensors to the human body, MMASD generates skeleton data by means of pre-trained deep neural networks. This is because children with autism have limited tolerance for external stimuli, and the presence of sensors on their bodies may cause them to become anxious, agitated, or exhibit challenging behaviors. Consequently, the skeleton data’s reliability in MMASD depends on the performance of the underlying pose detectors.
In addition, children with autism can exhibit varying motor functions, resulting in different intensity levels and completion rates for the same activity. Clinician evaluation results can offer additional guidance in determining and classifying activities for each individual. Nevertheless, there is a need to explore ways to incorporate tabulated clinician evaluation results with other movement-related features to generate comprehensive representations.
6. Conclusion
Autism research has been greatly facilitated by machine learning techniques, which offer cost-effective and accurate ways to analyze various aspects of children’s behavior and development. However, the lack of open-access datasets has posed challenges to conducting fair comparisons and promoting sound research practices in the field of autism research. To address this issue, we have proposed MMASD, a publicly accessible multimodal dataset that features diverse scenes collected from therapy interventions for children with autism. Our dataset includes multimodal features such as 2D & 3D skeleton data, optical flow, demographic data, and ADOS rating, offering a confidential data-sharing approach that can maintain critical motion information. MMASD distinguishes itself from existing works by utilizing privacy-preserving multimodal features to provide comprehensive representations of full-body movements across a range of diverse themes. Moreover, each scene in our dataset depicts the same activity performed by a child and one or more therapists, providing a valuable template for comparing typically developing individuals with children with ASD.
There are several directions for future work that are worth exploring. Firstly, further research can be conducted to develop and compare machine learning models on the MMASD dataset for various tasks, such as action quality assessment (Li et al., 2021a), interpersonal synchrony estimation (Li et al., 2022a, b), and cognitive status tracking. This can help establish benchmark performance and identify state-of-the-art methods for analyzing the full-body movements of children during therapeutic interventions. In addition, new approaches can be investigated to overcome pose detection failures in MMASD. For example, by introducing pose uncertainty or attention mechanism to assign higher weights to more reliable body joints. Furthermore, the MMASD dataset can be expanded in several aspects. To elaborate, further intervention scenarios could be included along with more features such as mutual gaze (Guo et al., 2023) and additional annotations like a synchrony score. Finally, efforts can be made to dataset augmentation via existing benchmarks not limited to the autism domain by matching samples with similar motion features (Pandey et al., 2020), which can significantly expand the scale of the autism dataset.